OpenStudy (anonymous):

Let f(x)=7x+2 and g(x)=2x^2-x-1. Compute: 5f(-1)-3g(-2) Would the f(x) and g(x) formula fit in between the letter and number? How would I be able to do this step by step?

OpenStudy (astrophysics):

f(-1) just means you plug in -1 wherever there is an x in your function, so f(-1) = 7(-1)+2

OpenStudy (anonymous):

So do it just like how you did the other problem, except we will multiply f(-1) by 5, and g(-2) by -3 .

OpenStudy (astrophysics):

and yes you're multiplying them by the function

OpenStudy (anonymous):

So the 5f and 3g are the final steps?

OpenStudy (anonymous):

Yes . Just remember it's -3 .

OpenStudy (jhannybean):

\[\color{red}{f(-1)=7(-1)+2} ~,~~ \color{blue}{g(-2)=2(-2)^2-(-2)-1}\]\[5(\color{red}{f(-1)})-3(\color{blue}{g(-2)})\]

OpenStudy (anonymous):

I had -4 as my answer?

OpenStudy (jhannybean):

That is not what I am getting.

OpenStudy (anonymous):

I don't get that

OpenStudy (anonymous):

hmm.. I am truly am bad at math. I guess I'll try to recalculate all that ><

OpenStudy (jhannybean):

First solve for f(-1), highlighted in red. Then do the same for g(-2) highlighted in blue

OpenStudy (astrophysics):

Oh wait multiply 9 by 3

OpenStudy (jhannybean):

Take your highlighted answers and just plug them into 5( ) - 3( )

OpenStudy (astrophysics):

Yes that should be -5 :P

OpenStudy (triciaal):

@Astrophysics 7(-1) + 2 = -5 not 5

OpenStudy (anonymous):

-2 ><

OpenStudy (astrophysics):

You can check the arithmetic

OpenStudy (triciaal):

@shawmoes not -9 should be 9

OpenStudy (astrophysics):

Yeah, -2 sounds good :)

OpenStudy (anonymous):

@triciaal Ahh yes, slip of the keyboard.

OpenStudy (astrophysics):

Uhggg

OpenStudy (astrophysics):

No not -2

OpenStudy (astrophysics):

f(-1) = 7(-1)+2 = -5 5*-5 = -25 g(-2) = 2(-2)^2 -(-2)-1 = 9 3(9) = 27 -25-27 =...

OpenStudy (astrophysics):

Damn signs haha!

OpenStudy (anonymous):

oh damn. Wait -52 0-0

OpenStudy (triciaal):

@reinatrumpet26 the process is correct.

OpenStudy (astrophysics):

The problem of doing problems without latex, screws everything up :P

OpenStudy (anonymous):

The problem is too many answerers , and then looking at their work and making a mistake because of their mistake ahaha.

OpenStudy (jhannybean):

Ok. -25-27 = - (25 + 27) = ? This is seriously not another 9 + 10 = 21 thing. Come on people

OpenStudy (astrophysics):

Yeah, I kept scrolling up and realizing there were arithmetic mistakes, what ever though, long as you understand the process, and yes I agree, too many people answering at the same time can get frustrating

OpenStudy (anonymous):

true I agree :) But its great to have more than one mind into thinking. Thanks for your help guys! Ill stick with my answer ^_^

OpenStudy (anonymous):

@Jhannybean Get off your high horse, we all make mistakes.

OpenStudy (jhannybean):

Wasn't sure if it was algebraic mistakes or distributive mistakes.